# Why are Sensitive Functions Hard for Transformers?

### v3

@@ Abstract
[…]
@@ 2 Model of Transformers
$y_{i}^{(k)}:=LayerNorm\left(f^{MLP}(Y_{i}^{(k)})\right)$ (4)
@@ 3 Average Sensitivity
$s(x,f):=\frac{1}{4}\sum_{i}|f(x)-f(x^{{}^{\oplus i}})|^{2}$ (6)
If $f$ maps to {±$\{\pm 1\}$, then $s(x,f)$ is the number of Hamming neighbors of $x$ on which $f$ flips. The average sensitivity for inputs of length $n$ is
@@
We explicitly define average sensitivity relative to input length $n$, as we will investigate the behavior of transformers performing a single function $f$ across varying input lengths.
@@
MAJORITY maps $x\in\{\pm 1\}^{n}$ to 1 if #$\#\{i:x_{i}=1\}>n/2$, and $-1$ else. Transformers show good length generalization (Merrill et al., 2022; Zhou et al., 2023). It is known that $as_{n}(f)\sim\sqrt{n}$ (O’Donnell, 2014). However, $s(f,x)=n$ whenever the ones and zeros in $x$ are almost fully balanced.
@@
MEAN maps $x\mapsto\frac{1}{n}\sum_{i=1}^{n}x_{i}$. We have $as_{n}(f)\sim\frac{1}{n^{2}}$.
@@ 6 Implications
For PARITY, $L_{\rho,n}$ converges to $1$ for large $d$, showing that arbitrarily small perturbations to a transformer computing PARITY will lead to a high loss for sufficiently long inputs. Previously, Chiang and Cholak (2022) noted that, for their hand-constructed transformer representing PARITY, small changes to a specific parameter led to a large increase in loss, suggesting that this made the solution impossible to reach with SGD. Theorem 6 shows that this phenomenon is unavoidable for any transformer representing a high-sensitivity function. For functions with sublinear average sensitivity, no such phenomenon is predicted.
@@
In the realm of multiplication over finite monoids, foundational to regular languages and the simulation of finite automata (e.g. Liu et al., 2023; Delétang et al., 2023; Angluin et al., 2023), average sensitivity creates a dichotomy: If $f$ indicates $n$-fold multiplication over a finite monoid, then $as_{n}(f)=O(1)$ if the monoid is aperiodic and $as_{n}(f)=\Theta(n)$ else. The theory thus predicts that transformer shortcuts to automata (Liu et al., 2023) will be brittle for many non-aperiodic automata.
PARITY can be solved well with a scratchpad (Anil et al., 2022; Liu et al., 2023). Existing theoretical accounts of the benefit of intermediate steps for transformers’ expressive capacity (e.g. Merrill and Sabharwal, 2023a; Feng et al., 2023) do not account for the benefit of intermediate steps for PARITY-like problems, as the upper bounds on transformer abilities used in these studies do not account for the hardness of learning to compute PARITY in a single step (Fact 1). The concept of average sensitivity provides a simple explanation. Formally, we can consider the problem of simulating a finite automaton with state set $\mathcal{X}$ either translating to the final state ${t}_{n}$ in one go (standard), or to autoregressively translate it into a sequence of states ${t}_{1},...,{t}_{n}$ (scratchpad). Then (proof in Appendix D):
@@ 7.1 Setup
We conducted experiments to test the predictions made by our theory, specifically assessing predictions regarding loss landscapes and sharpness of the minima. In all experiments, we use the transformer encoder architecture, using the default implementation in PyTorch (Paszke et al., 2019). Each model is trained to fit a function $f$ on a specific sequence length $n$. Each training input $x$ is generated uniformly from {±$\{\pm 1\}^{n}$, and each input bit, treated as a separate token, is embedded using learned token and positional encodings.
@@
    LayerNorm Blowup. This metric is computed by computing the maximum normalization factor (5) across the entire layer in each application of layer norm, and take the product over all applications of layer norm. This essentially corresponds to $\operatorname{Blowup}$ [^3] In the theory, Blowup has an additional 1+… in each factor for technical reasons. This difference is immaterial, as we are interested in situations where $Blowup=\omega(n)$.
@@ Acknowledgments
We thank David Chiang, Paul Lintilhac, and Yuval Pinter for spotting various errors in a previous version.
@@ Appendix A Simple Constructions for PARITY
A bit string is odd if and only if (normal-$\dagger$) for the (strict) majority of positions $i$ holding a 1, it holds that strictly more than 1/2 of 1s appear no later than at $i$.
@@ B.1 Bounding the Sensitivity of the Attention Heads
where (*)(*)( * ) holds with overwhelming probability over the choice of $x$, say, $1-\frac{1}{n^{{\delta}}}$ (multiplicative Chernoff bound for binomial random variables), for any individual $q$, and $1-\frac{1}{n^{{\delta-1}}}$ for the whole expression by a union bound over all $q$. We further used $\sum_{i=1}^{n}\frac{1}{i}\leq\log{n}+1$.
@@
where again (*)(*)( * ) holds with probability $1-\frac{1}{n^{\delta-1}}$ by a Chernoff bound for each $j$, and a union bound across all $j$. Importantly, the third inequality used the fact that $\gamma_{iw}\geq\gamma_{ij}$ when $w<j$. So overall
@@
[…]
@@ B.3 Layerwise Bounds on Sensitivity
$\boxed{\operatorname{I}[y_{j}^{(1)}]\leq 2\left(\max_{w}\|Y_{w}^{(1)}\|_{2}\right)\cdot N_{j}^{(1)}(x)\cdot N_{j}^{(1)}(x^{\oplus i})\cdot L_{f^{MLP}}\cdot\operatorname{I}[Y_{j}^{(1)}]}$ (50)
@@ Appendix C Theorem 6
$\geq$ $\mathbb{E}_{\Delta}\frac{(1-\frac{1}{\alpha})as_{n}(T_{\theta})}{(n-\frac{as_{n}(T_{\theta})}{\alpha\|T_{\theta}(x)\|_{2}^{2}})}+\|T_{\theta}\|_{2}^{2}\alpha\mathbb{E}_{\Delta}\frac{as_{n}(T_{\theta+\Delta})}{{as_{n}(T_{\theta})}}-2\sqrt{\frac{(1-\frac{1}{\alpha})as_{n}(T_{\theta})}{(n-\frac{as_{n}(T_{\theta})}{\alpha\|T_{\theta}(x)\|_{2}^{2}})}}\|T_{\theta}\|_{2}\mathbb{E}_{\Delta}\sqrt{\alpha\frac{as_{n}(T_{\theta+\Delta})}{{as_{n}(T_{\theta})}}}$
@@
where the last step used Jensen’s inequality. By the reasoning above, the factor 2 in the first term can be eliminated when the output is ±$\pm 1$. In order to lower-bound this, we want to show the following:

### v4

@@ Abstract
Michael Hahn, Mark Rofin Saarland Informatics Campus Saarland University, Saarbrücken, Germany {mhahn, mrofin}@lst.uni-saarland.de
@@ 2 Model of Transformers
$y_{i}^{(k)}:=LayerNorm\left(Y_{i}^{(k)}\right)$ (4)
@@ 3 Average Sensitivity
$s(x,f):=\frac{1}{4}\sum_{i=1}^{n}|f(x)-f(x^{{}^{\oplus i}})|^{2}$ (6)
where $x^{{}^{\oplus i}}$ denotes the bitstring $x$ with the $i$-th bit flipped. The average sensitivity for inputs of length $n$ is
@@
If $f$ maps to {±$\{\pm 1\}$, then $s(x,f)$ is the number of Hamming neighbors of $x$ on which $f$ flips. This definition of $as_{n}(f)$ corresponds to the “total influence” from O’Donnell (2014, Def. 2.27). We explicitly define average sensitivity relative to input length $n$, as we will investigate the behavior of transformers performing a single function $f$ across varying input lengths. The use of squared distances, rather than simple absolute distances, ensures that results about $as_{n}(f)$ transfer to results about degree profiles (Abbe et al., 2023), which we will later investigate (Eq. 17).
@@
MAJORITY maps $x\in\{\pm 1\}^{n}$ to 1 if #$\#\{i:x_{i}=1\}>n/2$, and $-1$ else. Transformers show good length generalization (Merrill et al., 2022; Zhou et al., 2023). It is known that $as_{n}(f)=\Theta(\sqrt{n})$ (O’Donnell, 2014, Ex. 2.22). However, $s(f,x)=n$ whenever the ones and zeros in $x$ are almost fully balanced.
@@
MEAN maps $x\mapsto\frac{1}{n}\sum_{i=1}^{n}x_{i}\in[-1,1]$. We have $as_{n}(f)=\frac{1}{n}$.
@@ 5 Sensitive Transformers are Brittle
While we stated Theorem 6 for an individual transformer, the same statement holds for families of transformers where weights may depend on $n$, as long as $C$ remains bounded. A consequence is that scaling attention logits with $\log n$ (used to represent sparse functions by Edelman et al. (2022); Chiang and Cholak (2022)) will, at least in the input layer, not mitigate the difficulty of sensitive functions.
@@ 6 Implications
For PARITY, $L_{\rho,n}$ converges to $1$ for large $d$, showing that arbitrarily small perturbations to a transformer computing PARITY will lead to a high loss for sufficiently long inputs. Previously, Chiang and Cholak (2022) noted that, for their hand-constructed transformer representing PARITY, small changes to a specific parameter led to a large increase in loss, suggesting that this made the solution impossible to reach with SGD. Theorem 6 shows that this phenomenon is unavoidable for any transformer representing a high-sensitivity function. For functions with sublinear average sensitivity, Theorem 6 entails no nontrivial lower bound on sharpness, and no such phenomenon is predicted.
@@
PARITY can be solved well with a scratchpad (Anil et al., 2022; Liu et al., 2023). Existing theoretical accounts of the benefit of intermediate steps for transformers’ expressive capacity (e.g. Merrill and Sabharwal, 2023a; Feng et al., 2023) do not account for the benefit of intermediate steps for PARITY-like problems: While the theoretical models of transformer expressivenes used in these studies do predict versions with intermediate steps to be easy, they do not predict that computing PARITY in a single step would be hard, due to Fact 1. The concept of average sensitivity provides a simple explanation for the benefit of intermediate steps. Formally, we can consider the problem of simulating a finite automaton with state set $\mathcal{X}$ either translating to the final state ${t}_{n}$ in one go (standard), or to autoregressively translate it into a sequence of states ${t}_{1},...,{t}_{n}$ (scratchpad). Then (proof in Appendix D):
@@ 7.1 Setup
We conducted experiments to test the predictions made by our theory, specifically assessing predictions regarding loss landscapes and sharpness of the minima. In all experiments, we use the transformer encoder architecture, using the default implementation in PyTorch (Paszke et al., 2019). Each model is trained to fit a function $f$ on a specific sequence length $n$. Each training input $x$ is generated uniformly from {±$\{\pm 1\}^{n}$, and each input bit, treated as a separate token, is embedded using learned token and positional encodings. [^3] Code is available at https://github.com/lacoco-lab/sensitivity-hardness.
@@
    LayerNorm Blowup. This metric is computed by computing the maximum normalization factor (5) across the entire layer in each application of layer norm, and take the product over all applications of layer norm. This essentially corresponds to $\operatorname{Blowup}$ [^4] In the theory, Blowup has an additional 1+… in each factor for technical reasons. This difference is immaterial, as we are interested in situations where $Blowup=\omega(1)$.
@@ Acknowledgments
We thank Lena Strobl, Dana Angluin, and David Chiang for useful discussion. We thank David Chiang, Paul Lintilhac, and Yuval Pinter for spotting various errors in a previous version, and the anonymous ARR reviewers for useful feedback.
@@ Appendix A Simple Constructions for PARITY
While it is well-known that PARITY can be expressed using uniform circuits with majority gates (haastad1986computational), it may not be obvious just how simple the construction can be, and how it can easily embed into formalisms like RASP. In order to highlight the existence of very simple constructions for PARITY in various formalisms, and to make discussion here self-contained, we provide a simple proof. The key is the following insight:
@@
A bit string is odd if and only if ($\dagger$) for the (strict) majority of positions $i$ holding a 1, it holds that strictly more than 1/2 of 1s appear no later than at $i$.
@@ B.1 Bounding the Sensitivity of the Attention Heads
where ()(*)( ∗ ) holds with overwhelming probability over the choice of $x$, say, $1-\frac{1}{n^{{\delta}}}$ (multiplicative Chernoff bound for binomial random variables), for any individual $q$, and $1-\frac{1}{n^{{\delta-1}}}$ for the whole expression by a union bound over all $q$. We further used $\sum_{i=1}^{n}\frac{1}{i}\leq\log{n}+1$.
@@
where again ()(*)( ∗ ) holds with probability $1-\frac{1}{n^{\delta-1}}$ by a Chernoff bound for each $j$, and a union bound across all $j$. Importantly, the third inequality used the fact that $\gamma_{iw}\geq\gamma_{ij}$ when $w<j$. So overall
@@
$\widehat{a}_{i,j}^{(1,h)}=\begin{cases}\frac{n}{n+\frac{n-1}{n}}=\frac{1}{1+1/n-1/n^{2}}\geq 1-\frac{1}{n}&\text{if }j=q\\ \frac{1/n}{n+\frac{n-1}{n}}=\frac{1}{n^{2}+n-1}\leq\frac{1}{n^{2}}&\text{else}\end{cases}$ (39)
@@ B.3 Layerwise Bounds on Sensitivity
$\boxed{\operatorname{I}_{i}[y_{j}^{(1)}]\leq 2\left(\max_{w}\|Y_{w}^{(1)}\|_{2}\right)\cdot N_{j}^{(1)}(x)\cdot N_{j}^{(1)}(x^{\oplus i})\cdot L_{f^{MLP}}\cdot\operatorname{I}_{i}[Y_{j}^{(1)}]}$ (50)
@@ Appendix C Theorem 6
$\geq$ $\mathbb{E}_{\Delta}\underbrace{\frac{(1-\frac{1}{\alpha})as_{n}(T_{\theta})}{(n-\frac{as_{n}(T_{\theta})}{\alpha\|T_{\theta}(x)\|_{2}^{2}})}}_{\text{independent of $\Delta$}}+\underbrace{\|T_{\theta}\|_{2}^{2}\alpha\mathbb{E}_{\Delta}\frac{as_{n}(T_{\theta+\Delta})}{{as_{n}(T_{\theta})}}}_{\geq 0}-2\sqrt{\frac{(1-\frac{1}{\alpha})as_{n}(T_{\theta})}{(n-\frac{as_{n}(T_{\theta})}{\alpha\|T_{\theta}(x)\|_{2}^{2}})}}\|T_{\theta}\|_{2}\mathbb{E}_{\Delta}\sqrt{\alpha\frac{as_{n}(T_{\theta+\Delta})}{{as_{n}(T_{\theta})}}}$
$\geq$ $\frac{as_{n}(T_{\theta})}{2n-\frac{as_{n}(T_{\theta})}{\|T_{\theta}(x)\|_{2}^{2}}}-2\sqrt{\frac{1}{(n-\frac{1}{2\|T_{\theta}(x)\|_{2}^{2}})}}\|T_{\theta}\|_{2}\mathbb{E}_{\Delta}\sqrt{as_{n}(T_{\theta+\Delta})}$
$\geq$ $\frac{as_{n}(T_{\theta})}{2n-\frac{as_{n}(T_{\theta})}{\|T_{\theta}(x)\|_{2}^{2}}}-2\mathbb{E}_{\Delta}\sqrt{\frac{as_{n}(T_{\theta+\Delta})}{(n-\frac{1}{2\|T_{\theta}(x)\|_{2}^{2}})}}$
$\geq^{(\dagger)}$ $\frac{as_{n}(T_{\theta})}{2n-\frac{as_{n}(T_{\theta})}{\|T_{\theta}(x)\|_{2}^{2}}}-4\mathbb{E}_{\Delta}\sqrt{\frac{as_{n}(T_{\theta+\Delta})}{n}}$
$=$ $\frac{as_{n}(T_{\theta})}{2n-\frac{as_{n}(T_{\theta})}{\|T_{\theta}(x)\|_{2}^{2}}}-4\mathbb{E}_{\Delta}\sqrt{\frac{\mathbb{E}_{x}s_{n}(T_{\theta+\Delta},x)}{n}}$
@@
where ()($\dagger$)( † ) uses $(n-\frac{1}{2\|T_{\theta}(x)\|_{2}^{2}})\geq n-1/2\geq n/4$, and the final step used Jensen’s inequality applied to the concave function $\sqrt{\cdot}$. By the reasoning above, in Eq. (67), the factor 2 in the first term can be eliminated when the output is ±$\pm 1$. In order to lower-bound (68), we want to show the following: